OpenStudy (anonymous):

y=-2x^2+5x-1 x=0 y=0

OpenStudy (solomonzelman):

What do you need to do? Find the y-intercept and then the x-intercepts ?

OpenStudy (anonymous):

y intercept

OpenStudy (solomonzelman):

to find the y-intercept, set x=0 and solve for y. the first 2 terms 2x^2 and 5x if you make x=0 are going to turn into zero, so what's left ?

OpenStudy (anonymous):

-1

OpenStudy (solomonzelman):

Yes, there we go !

OpenStudy (anonymous):

how do we solve for x

OpenStudy (solomonzelman):

How would you solve for x-intercepts ? set y=0 and solve for x. Like this \[y=-2x^2+5x-1 \]\[-2x^2+5x-1 =0\] multiply everything times -1 to make it look easier :)\[2x^2-5x+1 =0\] (you can't factor since the discriminant isn't a perfect square, but, you can use the quadratic formula) You want to do the quadratic formula, or should I start you off first ?

OpenStudy (anonymous):

could you start it off not sure how to